How much water (density = 1.00 X 103 kg/m3 ) must be displaced to float a cubic block of wood (density 655 kg/m3) that is 15.0 cm on a side?

1
Expert's answer
2021-09-24T09:25:57-0400

"a = 15cm = 0.15m"

"\rho_w = 1*10^3\frac{kg}{m^3}"

"\rho _c= 655\frac{kg}{m^3}"

"g = 9.8 \frac{kg}{m^3}"

"\text{Forces acting on a wooden cube:}"

"F_A = \rho_wgV_1"

"V_1 - \text{The volume of a part of the cube immersed in}"

"\text{water and also the volume of displaced water}"

"F_g = m_cg"

"m_c- \text{mass of a wooden cube}"

"m_c = V_c\rho_c = a^3\rho_c"

"F_g = F_A"

"a^3\rho_c g = \rho_wgV_1"

"V_1 = \frac{a^3\rho_c} {\rho_wg}= \frac{0.15^3*665}{10^3}=2.24*10^{-3}m^3"

"m_1 -\text{mass of displaced water}"

"m_1 = \rho_wV_1= 2.24kg"

"\text{Answer:}"

"2.24kg-\text{mass of displaced water}"

"2.24*10^{-3}m^3 \text{-volume of displaced water}"